Well my teacher wants us to use the Distributive Property and combine like terms in these questions, but I'm stuck on three questions, Please Help!
1. 4xsquared + 12y + 3xsquared - 2y =
2. 3(x + 5) + 4 =
3. (x + 3) + 6 =
I need some understanding.

Answer by guapa(62)

No 1: 4x² +12y +3x² -2y
Your like terms are 4x²+3x²=(4+3)x² =7x² and 12y-2y=(12-2)y =10y. You just need to add or substract the coefficient of the terms with the same variable.
So, the result to No 1 = 7x²+10y
No 2: 3(x+5)+4
The distributive property says: a(b+c)=ab+ac. So in this case we get:
3(x+5)+4= 3x+15+4 Now combine like terms
3x+19 This is your final answer
No 3: (x+3)+6
(x+3)means actually 1(x+3). But we don't write this.
So, simply proceed as follows:
(x+3)+4 = x+3+4 = x+7
Hope this helps